On August 1, 2011, Bonnie purchased $20,000 of Huber Co.'s 14%, 13-year bonds at face value. Huber co. has paid the semiannual interest due on the bonds regularly. On August 1, 2019, market rates of interest had fallen to 12% and Bonnie is considering selling the bonds. Calculate the market value of Bonnie's bonds on August 1, 2019.
Answer #1

14% Bond of Huber Co Purchased by bonnie at August 1 2011 = $20000

On August 1 2019 the Interest Rate of Bond decline from 14% to 12%

Therefore % of decline in Interest Rate is = 14* (14-12)/100 = 0.28%

General Principle - When interest rate of Bond decreases the bond price increases. The Bond price and the interest rate are inversly proportional.The interest rates will lower for a high yielding bond.

Therefore the Market Price of Bond will increase by 0.28%

The market price of Bond = 20000* (100+0.28)/100 = $20056